Installation & Operation Status.


  • 10m Mode Cleaner (97.4-97.12)
    Sidebands(15.25MHz) transmittion experiment was successfully finished. Residual excess noise is smaller than 20mW shot noise.
  • One Arm (97.8-98.4)
    300m FP cavity was successfully locked. By using a wave front sensing method, the cavity is very stable.
  • Laser & Mode Cleaner (98.1-99.1)
    3.5W output after Mode Cleaner has been realized! Transmittion rate is 54%.
  • Recombine I (98.5-99.1)
    Operation of Fabry-Perot Michelson Interferometer(FPMI) succeeded(18 Sep.). Five hours' data was taken.
  • Connection (99.2)
    Laser, MC and FPMI were connected.
  • Recombine II (99.3-00.8)
    Operation of FPMI with 10W Laser succeeded using pre-modulation method(22 May, 1999). Best sensitivity is h = 5 x 10-21 /sqrt(Hz). This is world record. (14 Aug, 2000).
    167 hour data was taken in Data Taking 4(21 Aug. -4 Sep., 2000).
  • Recombine III (00.9-01.10)
    Active isolator was installed. Mirror suspensions and a matching telescope were modified.
  • Full-time Observation (01.8-01.9)
    Total 1038 hours data was taken. Duty cycle was 86.5%.
  • Recycling (01.11-05.8)
    Best sensitivity is h = 2 x 10-21 /sqrt(Hz). Data taking 7, 8, and 9 were held.
  • Recycled Michelson Experiments (04.1-04.12)
    In order to investigate the noise behavior, experiments were done using the recycled Michelson interferometer without 300m arms.
  • SAS (05.9-08.3)
    The sensitivity was improved at the all frequency range due to the seismic attenuation system (SAS).
  • RSE (08.4-)
    Now going.